# HD-OOD3D: Supervised and Unsupervised Out-of-Distribution object detection in LiDAR data

### v2

@@ Abstract
Autonomous systems rely on accurate 3D object detection from LiDAR data, yet most detectors are limited to a predefined set of known classes, making them vulnerable to unexpected out-of-distribution (OOD) objects. In this work, we present HD-OOD3D, a novel two-stage method for detecting unknown objects. We demonstrate the superiority of two-stage approaches over single-stage methods, achieving more robust detection of unknown objects while addressing key challenges in the evaluation protocol. Furthermore, we conduct an in-depth analysis of the standard evaluation protocol for OOD detection, revealing the critical impact of hyperparameter choices. To address the challenge of scaling the learning of unknown objects, we explore unsupervised training strategies to generate pseudo-labels for unknowns. Among the different approaches evaluated, our experiments show that top-5 auto-labelling offers more promising performance compared to simple resizing techniques.
@@ I INTRODUCTION
LiDAR sensors, offering superior depth estimation and high-resolution 3D point cloud data, have become essential for autonomous vehicles, environmental mapping, and infrastructure monitoring, driving advances in 3D object detection as their availability and affordability increase. Despite significant advancements, current neural network-based 3D object detection models still encounter substantial challenges when faced with ”unknown” classes, objects that were not part of the training dataset.
In real-world scenarios, encountering objects that lie outside the training distribution—referred to as out-of-distribution (OOD) objects—is inevitable and can pose significant safety risks. This issue is further exacerbated by the fact that the annotation of diverse object classes is both labor-intensive and costly, often leading to datasets that are not fully representative of real-world complexity.
@@
  • Evaluating generation methods of additional data for training an OOD-aware 3D object detector.
@@ II-B Open-Set and Out-Of-Distribution Detection
In 3D object detection, handling OOD scenarios has also seen substantial progress. Cen et al. [1] explored methods for detecting and managing unknown objects within 3D environments. However, their evaluations are based on UDI, a private dataset with only 600 unknown test objects, and KITTI, which we will show is insufficient. UFO [14] introduced novel approaches for identifying unknown objects in point clouds, focusing on accurately segmenting foreground objects, but they also focus on the limited KITTI dataset. Huang et al. [15] explored different instance injection methods to test OOD baselines. The variance in their results underlines the difficulty in identifying the OOD objects they insert from different sources, which have different representations from the KITTI dataset in which they are injected. Lastly, Kösel et al. [16] offered an evaluation of existing OOD techniques in the context of LiDAR data, identifying key challenges and proposing enhancements. However, their work lacks specifications on their evaluation protocol and evaluated objects, such as the recall/hit percentage of unknowns, and only test resizing unknown generation on nuScenes, which has certain limitations as we will see in Section IV. We aim to improve the evaluation methodology behind their work as well as the performance through a rigorous benchmark of unknown generation. We find that using a focal loss instead of a binary cross-entropy loss yields better results. Furthermore, their resizing unknown generation method produces lesser results than top-K labelling, suggesting that using pseudo-labels from real OOD data can lead to better performance than reshaping known objects.
@@ III A new method for 3D OOD Detection
While traditional open-set detectors add an extra “unknown” class, we instead view the challenge as an OOD prediction problem. Closed-set detectors are optimized for a fixed set of “core” classes, but an OOD framework allows the model to flag deviations from the learned distribution without sacrificing known-class performance. In this section, we compare methods that compute “unknown” object scores directly from the detector’s outputs (single-stage methods) with those that add an extra trainable module (two-stage methods). We also introduce a novel input representation for two-stage methods.
III-A Datasets for OOD Evaluation
III-A1 The limits of widely used datasets
@@
nuScenes [20], while one of the most comprehensive datasets, presents specific challenges critical in the context of OOD detection. Firstly, the predefined dataset splits and evaluation metrics in nuScenes might not be sufficient for evaluating OOD performance. For instance, only one out of four of the dataset’s scans in its evaluation set contains unknown objects as defined by [16], thus limiting the evaluation interpretation. nuScenes also suffers from its point cloud sparsity due to its 32-beam LiDAR sensor, which can only be solved by aggregating scans and creating a trailing point cloud for moving objects, hindering detection. These findings underscore the need to use a more diverse and dense dataset to effectively tackle the OOD problem in 3D object detection.
III-A2 Argoverse 2
@@
Dataset Hits % (Open) Hits % (Closed)
@@
For Argoverse 2, we choose the rarest 5% of classes as our unknown classes, denoted as set $U$. This approach ensures that the chosen unknown classes are genuinely underrepresented in the training data, providing a challenging test for the OOD detection capabilities of our models. This selection method helps simulate real-world scenarios where we assume that the most common objects will have been labeled beforehand while certain rare objects are infrequently encountered, making it a valuable benchmark for assessing the robustness of 3D object detection systems. The unknown classes $U$, representing $28,271$ instances in the validation set, are:
[…]
@@
[…]
The closed ”known” classes, denoted as set $C$, include more common objects, presumed to have been labeled beforehand. These classes, representing $1,391,546$ instances in the validation set, are:
[…]
@@
[…]
Methods to compute ”unknown” object scores can be generally separated into two categories: single-stage methods, which directly use the outputs of a detector model and apply some thresholding or entropy computation, and two-stage methods, which add additional trainable parameters on top of the base detector. This latter approach necessitates a retraining phase.
III-B Single-Stage Methods
@@
In our evaluation of OOD detection for 3D object detection, we selected several single-stage methods to ensure a comprehensive assessment, i.e. methods that directly use the output logits of the detection model. The chosen methods include ODIN [10], MaxLogit [11], Energy Thresholding [12], MC Dropout inspired by [4], and the Default Score.
III-C Two-Stage Methods
We introduce our proposed method High-Dimensional OOD 3D Object Detection (HD-OOD3D), which integrates an additional Multi-Layer Perceptron (MLP) module after the base detector. This module is specifically trained to output a binary classification confidence score or an ”unknown” confidence score: 0 for known objects and 1 for unknown objects. We illustrate the training pipeline of HD-OOD3D in Figure 1. This two-stage solution employs the MLP to process initial detections, decomposed into embeddings $e_{i}$ extracted at the feature map locations of predicted bounding boxes, box predictions $b_{i}$, and prediction logits $p_{i}$ and outputs an ”out-of-distribution” score $p_{\text{OOD}}$. During inference, if $p_{\text{OOD}}$ is larger than a given threshold $\text{OOD}_{\text{thresh}}$, the object is considered an OOD object. Unlike MMOOD3D[16], which uses lower-resolution input representation and generates unknowns using a resizing method, we use different input representations and unknown generation methods, the benefits of which we demonstrate in Section III-E and Section V-B respectively. In the first stage, we train a CenterPoint [7] object detector on the Argoverse 2 dataset [21] with classical training regimes [16, 7]. This model obtained a mAP score of 35.2 on the known classes. We freeze this first stage model then add a 3-layer MLP where we divide the dimensionality by two between every layer, followed by a sigmoid activation function. We train it for 5 epochs, with a polynomial learning rate using a power of $3$, with an initial learning rate of $1\mathrm{e}{-3}$.
First, using a methodology we will refer to as ”supervised,” we will leverage the labels of the unknown objects $U$ present in the Argoverse 2 dataset. This supervised approach allows for a more rigorous comparison of the different methods: one-stage, two-stage, and our HD-OOD3D method against the state of the art. In the second part, Section V, we will explore the ”unsupervised” methodology, where we assume we do not have access to the labels of the unknown objects $U$ in the Argoverse 2 dataset. This will enable us to compare various pseudo-label generation methods for unknown objects.
III-D Results on unknown learning with single and two-stage methods
@@
We compared multiple single-stage methods for ODD detection on the Argoverse 2 dataset with our evaluation protocol of Section IV. We assess the OOD detector using several metrics: AUROC, which quantifies the overall likelihood that a positive case is ranked above a negative case; FPR-95, which indicates the false positive rate when the true positive rate is set at 95%, balancing sensitivity and specificity; and the AUPR metrics, AUPR-E and AUPR-S, which focus on the model’s precision-recall balance for unknown and known objects respectively. Importantly, AUPR-S represents the model performance on known objects. If a noticeable downgrade in this metric is noticed, such as under 99, then we consider the base detector’s performance on known objects to be hindered by the additional unknown object classifier module. Table II shows the results. All single-stage models fail to reach the performance of two-stage methods. While ODIN achieves the highest AUROC among single-stage approaches (85.9 on Argoverse 2), it still exhibits a high False Positive Rate at 95% recall (61.6%). Energy-based methods, Monte-Carlo Dropout, MaxLogit, and Default Score thresholding yield relatively weaker results, with AUROC values ranging from 65.8 to 76.6. In contrast, two-stage methods significantly outperform single-stage approaches. MMOOD3D reaches an AUROC of 88.6 with a lower FPR-95 of 43.2. Our proposed method, HD-OOD3D, achieves competitive results with an AUROC of 86.8 while maintaining a lower FPR-95 (51.1) compared to ODIN. Crucially, HD-OOD3D preserves the performance on known objects (AUPR-S = 99.7), whereas MMOOD3D suffers a substantial drop in AUPR-S to 70.5. This suggests that HD-OOD3D provides a better balance between detecting unknown objects and preserving the accuracy of the base detector.
@@
TABLE II: Comparison of AUROC, FPR-95, AUPR-S, AUPR-E metrics for single-stage and two-stage methods on Argoverse 2 in the supervised setup, where ground truth labels for unknown objects are used during training.
III-E Ablation study on HD-OOD3D
The two-stage methods use an MLP for OOD prediction. We first compare different inputs of the MLP using the labels of unknown objets as previously:
  • feature interpolation (Interp.): Instead of extracting the feature representation of the base detector’s boxes at the discrete grid value of the feature map, we instead interpolate these features based on the predicted center
  • using a focal loss instead of a binary cross-entropy loss (F. Loss)
  • using a $3\times 3$ max-pooling (Pool) of the feature map before extracting the feature vector at the corresponding feature map location of the predicted object center
@@ III-E Ablation study on HD-OOD3D
TABLE III: Comparison of AUROC, FPR-95, AUPR-S, AUPR-E metrics for different variations of an MLP two-stage method on Argoverse 2. The configuration represents that of [16] while the grey line is the one retained in all experiments in this work.
We show our experiments using the ground truth unknown objects for training in Table III. The results show that adding a 3x3 pooling layer and a focal loss was beneficial, compared to the work [16], which only uses a binary cross-entropy loss, no pooling, high-dimensional feature maps, and feature interpolation. Despite a higher FPR-95, this is the only method that combines a higher precision on known objects and near-perfect precision on closed objects. For instance, in the Argoverse 2 dataset, a $3\times 3$ maxpooling filter enabled a feature aggregation operation with a radius of 2.1 meters. It substantially improved OOD performance by ensuring a more robust aggregation of surrounding point cloud data.
@@ IV Analysis of the Evaluation Protocol for 3D OOD
Here, we examine how external factors and hyperparameter choices influence OOD detection performance because these elements directly affect both the quality and reliability of the detection metrics. For instance, the characteristics of the dataset—such as point cloud density, class diversity, and the presence of noise—can alter how many true and false positives are identified. Similarly, hyperparameters like the distance threshold for matching predictions, the detection score cutoff, and the criteria used for sorting predictions determine which detections are considered valid. Even small adjustments in these parameters can lead to significant shifts in performance metrics, affecting the reproducibility and fairness of the evaluation. Therefore, we examine and compare these choices.
@@ IV-A Metrics and Matching
       Compute Euclidean distance between the M predictions and the $N$ GTs. Sort by order of confidence the distances, the boxes, and ”unknown” scores $p_{\text{OOD}}$;
@@
                   if dist(pred, closest unmatched GT)$<$ $d_{\text{thresh}}$ then

### v3

@@ Abstract
Autonomous systems rely on accurate 3D object detection from LiDAR data, yet most detectors are limited to a predefined set of known classes, making them vulnerable to unexpected out-of-distribution (OOD) objects. In this work, we present HD-OOD3D, a novel two-stage method for detecting unknown objects. We demonstrate the superiority of two-stage approaches over single-stage methods, achieving more robust detection of unknown objects. Furthermore, we conduct an in-depth analysis of the standard evaluation protocol for OOD detection, revealing the critical impact of hyperparameter choices. To address the challenge of scaling the learning of unknown objects, we explore unsupervised training strategies to generate pseudo-labels for unknowns. Among the different approaches evaluated, our experiments show that top-K auto-labelling offers more promising performance compared to simple resizing techniques.
@@ I INTRODUCTION
LiDAR sensors, offering superior depth estimation and high-resolution 3D point cloud data, have become essential for autonomous vehicles, environmental mapping, and infrastructure monitoring, driving advances in 3D object detection as their availability and affordability increase. Despite significant advancements, current neural network-based 3D object detection models still encounter substantial challenges when faced with ”unknown” objects that were not part of the training dataset [1].
In real-world scenarios, encountering objects that lie outside the training distribution, referred to as out-of-distribution (OOD) objects, is inevitable and can pose significant safety risks. This issue is further exacerbated by the fact that the annotation of diverse object classes is both labor-intensive and costly, often leading to datasets that are not fully representative of real-world complexity.
@@
  • Evaluating generation methods of data for unsupervised training an OOD-aware 3D object detector.
@@ II-B Open-Set and Out-Of-Distribution Detection
In 3D object detection, handling OOD scenarios has also seen substantial progress. Cen et al. [1] explored methods for detecting and managing unknown objects within 3D environments. However, their evaluations are based on UDI, a private dataset with only 600 unknown test objects, and KITTI dataset [14]. UFO [15] introduced novel approaches for identifying unknown objects in point clouds, focusing on accurately segmenting foreground objects, but they also focus only on the KITTI dataset [14]. Huang et al. [16] explored different instance injection methods to test OOD baselines. The variance in their results underlines the difficulty in identifying the OOD objects they insert from different sources, which have different representations from the KITTI dataset in which they are injected. Lastly, Kösel et al. [17] offered an evaluation of existing OOD techniques in the context of LiDAR data, identifying key challenges and proposing enhancements. However, their work lacks specifications on their evaluation protocol and evaluated objects, such as the recall/hit percentage of unknowns, and simply resizes known objects to generate unknown ones and evaluates their method exclusively on nuScenes, which has certain limitations, as we will discuss in Section III-B1. We aim to improve the evaluation methodology behind their work as well as the performance through a rigorous benchmark of unknown generation. We find that using a focal loss instead of a binary cross-entropy loss yields better results. Furthermore, their resizing unknown generation method produces lesser results than top-K labelling, suggesting that using pseudo-labels from real OOD data can lead to better performance than reshaping known objects.
@@ III A new method for 3D OOD Detection
While traditional open-set detectors add an extra “unknown” class, we instead view the challenge as an OOD prediction problem. Closed-set detectors are optimized for a fixed set of “core” classes, but an OOD framework allows the model to flag deviations from the learned distribution without sacrificing known-class performance. Methods to compute ”unknown” object scores can be generally separated into two categories: single-stage methods, which directly use the outputs of a detector model and apply some thresholding or entropy computation, and two-stage methods, which add additional trainable parameters on top of the base detector. This latter approach necessitates a retraining phase.
In this section, we introduce our new approach for OOD detection and compare methods that compute “unknown” object scores directly from the detector’s outputs (single-stage methods) with those that add an extra trainable module (two-stage methods).
III-A Our approach
We introduce our proposed method High-Dimensional OOD 3D Object Detection (HD-OOD3D), which integrates an additional Multi-Layer Perceptron (MLP) module after the base detector. This module is specifically trained to output a binary classification confidence score or an ”unknown” confidence score: 0 for known objects and 1 for unknown objects. We illustrate the training pipeline of HD-OOD3D in Figure 1. This two-stage solution employs the MLP to process initial detections, decomposed into embeddings $e_{i}$ extracted at the feature map locations of predicted bounding boxes, box predictions $b_{i}$, and prediction logits $p_{i}$ and outputs an ”out-of-distribution” score $p_{\text{OOD}}$. During inference, if $p_{\text{OOD}}$ is larger than a given threshold $\text{OOD}_{\text{thresh}}$, the object is considered an OOD object.
The closest approach to our method is MMOOD3D [17], which is also a two-stage method. The main differences with MMOOD3D are:
  • feature interpolation (Interp.): instead of extracting the feature representation of the base detector’s boxes at the discrete grid value of the feature map, we interpolate these features based on the predicted center
  • a focal loss instead of a binary cross-entropy loss (F. Loss)
  • a $3\times 3$ max-pooling (Pool) of the feature map before extracting the feature vector at the corresponding feature map location of the predicted object center
Section III-D discusses the contribution of these modifications. Section V-B also presents a comparison between our approach to unsupervised learning of unknowns using top-K labeling and the MMOOD3D method [17], which constructs unknown objects by resizing known ones.
III-B Evaluation with supervised unknown objects
First, using a methodology we will refer to as ”supervised,” we will leverage the labels of the unknown objects in the chosen dataset. This supervised approach allows for a more rigorous comparison of the different methods: one-stage, two-stage, and our HD-OOD3D method. In the second part, Section V, we will explore the ”unsupervised” methodology, where we assume we do not have access to the labels of the unknown objects in the dataset. This will enable us to compare various pseudo-label generation methods for unknown objects.
III-B1 The limits of widely used datasets
@@
nuScenes [20] is one of the most used datasets in autonomous driving but presents specific challenges critical in the context of OOD detection. Firstly, the predefined dataset splits and evaluation metrics in nuScenes might not be sufficient for evaluating OOD performance. For instance, only one out of four of the dataset’s scans in its evaluation set contains unknown objects as defined by [17], thus limiting the evaluation interpretation. nuScenes also suffers from its point cloud sparsity due to its 32-beam LiDAR sensor, which can only be solved by aggregating scans and creating a trailing point cloud for moving objects, hindering detection. These findings underscore the need to use a more diverse and dense dataset to effectively tackle the OOD problem in 3D object detection.
III-B2 Argoverse 2 dataset for OOD evaluation
@@
Dataset Hits % Unknown objects Known objects
@@
For Argoverse 2, we choose the rarest 5% of classes as our unknown classes, denoted as set $U$. This approach ensures that the chosen unknown classes are genuinely underrepresented in the training data, providing a challenging test for the OOD detection capabilities of our models. This selection method helps simulate real-world scenarios where we assume that the most common objects will have been labeled beforehand while certain rare objects are infrequently encountered, making it a valuable benchmark for assessing the robustness of 3D object detection systems. The unknown classes $U$ (11 labels), representing $28,271$ instances in the validation set, are:
$U$ $=\{\text{Motorcyclist, School bus, Message board trailer, Truck cab,}$
@@
$\quad\text{sign, Wheeled rider, Wheelchair, Dog}\}$
The closed ”known” classes, denoted as set $C$ (15 labels), include more common objects, presumed to have been labeled beforehand. These classes, representing $1,391,546$ instances in the validation set, are:
$C$ $=\{\text{Regular vehicle, Pedestrian, Bollard, Construction cone,}$
@@
$\quad\text{Box truck, Large vehicle, Construction barrel, Vehicular trailer}\}$
III-B3 Implementation and metrics
In the first stage, we train a CenterPoint [7] object detector on the Argoverse 2 dataset [21] with classical training regimes [17, 7]. Our implementation is based on the OpenPCDet framework [22]. Specifically, CenterPoint is trained and evaluated in the $\left[-72.0,72.0\right]\mathrm{m}$ x-y range and uses a voxel size of $\left[0.1,0.1,0.2\right]\mathrm{m}$. We augment the data using random flipping along the x-axis, rotation around the z-axis, and random scaling of the point clouds in the $\left[0.95,1.05\right]$ range. We use the Adam optimizer with a OneCycle learning rate, initialized at $1\mathrm{e}{-3}$, a weight decay of $0.01$, momentum of $0.9$, and train for ten epochs, using a focal loss to compute the gradients.
This model obtained a mAP score of $35.2$ on the known classes (15 labels). We freeze this first stage model then add a 3-layer MLP where we divide the dimensionality by two between every layer ($512\rightarrow 256\rightarrow 128\rightarrow 2$), followed by a sigmoid activation function. We train it for five epochs, with a polynomial learning rate using a power of $3$, with an initial learning rate of $1\mathrm{e}{-3}$.
The proposed second-stage neural network adds only about 500k parameters and a few milliseconds of inference time, both negligible compared to the 30M parameters and latency (around 50 ms) of the base CenterPoint detector. This demonstrates that two-stage methods can provide OOD capability with minimal overhead, making HD-OOD3D feasible for real-world deployment in autonomous vehicles.
@@ III-B3 Implementation and metrics
We compared multiple single-stage methods for ODD detection and also the two-stage method MOOD3D [17] against our approach on the Argoverse 2 dataset with our evaluation protocol of Section IV. We selected several single-stage methods to ensure a comprehensive assessment, i.e. methods that directly use the output logits of the detection model. The chosen methods include ODIN [10], MaxLogit [11], Energy Thresholding [12], MC Dropout inspired by [4], and the Default Score.
@@
We assess the OOD detector using several metrics: AUROC, which quantifies the overall likelihood that a positive case is ranked above a negative case; FPR-95, which indicates the false positive rate when the true positive rate is set at 95%, balancing sensitivity and specificity; and the AUPR metrics, AUPR-E and AUPR-S, which focus on the model’s precision-recall balance for unknown and known objects respectively. Importantly, AUPR-S represents the model performance on known objects. If a noticeable downgrade in this metric is noticed, such as under 99, then we consider the base detector’s performance on known objects to be hindered by the additional unknown object classifier module. The threshold $\text{OOD}_{\text{thresh}}$ is dynamically tuned at evaluation to compute metrics like AUROC, FPR-95, and AUPR.
III-C Results
Table II shows the results. All single-stage models fail to reach the performance of two-stage methods. While ODIN achieves the highest AUROC among single-stage approaches (85.9 on Argoverse 2), it still exhibits a high False Positive Rate at 95% recall (61.6%). Energy-based methods, Monte-Carlo Dropout, MaxLogit, and Default Score thresholding yield relatively weaker results, with AUROC values ranging from 65.8 to 76.6. In contrast, two-stage methods significantly outperform single-stage approaches. MMOOD3D reaches an AUROC of 88.6 with a lower FPR-95 of 43.2. Our proposed method, HD-OOD3D, achieves competitive results with an AUROC of 86.8 while maintaining a lower FPR-95 (51.1) compared to ODIN. Crucially, HD-OOD3D preserves the performance on known objects (AUPR-S = 99.7), whereas MMOOD3D suffers a substantial drop in AUPR-S to 70.5. This suggests that HD-OOD3D provides a better balance between detecting unknown objects and preserving the accuracy of the base detector.
@@ III-C Results
TABLE II: Comparison of AUROC, FPR-95, AUPR-S, AUPR-E metrics for single-stage and two-stage methods on Argoverse 2 in the supervised setup, where ground truth labels for unknown objects are used during training. The yellow cells indicate methods with an AUPR-S score below 99, which means that the base detector has lost performance on the known classes.
III-D Ablation study on HD-OOD3D
In Table III, we compare different parts of our approach described in Section III-A.
@@ III-D Ablation study on HD-OOD3D
TABLE III: Comparison of AUROC, FPR-95, AUPR-S, AUPR-E metrics for different variations of an MLP two-stage method on Argoverse 2. The configuration represents method MMOOD3D [17] while the grey line is our approach HD-OOD3D. The yellow cells indicate methods with an AUPR-S score below 99, which means that the base detector has lost performance on the known classes.
The results show that adding a 3x3 pooling layer and a focal loss was beneficial, compared to the work [17], which only uses a binary cross-entropy loss, no pooling, high-dimensional feature maps, and feature interpolation. Despite a higher FPR-95, this is the only method that combines a higher precision on known objects and near-perfect precision on closed objects. For instance, in the Argoverse 2 dataset, a $3\times 3$ maxpooling filter enabled a feature aggregation operation with a radius of 2.1 meters. It substantially improved OOD performance by ensuring a more robust aggregation of surrounding point cloud data.
@@ IV Analysis of the Evaluation Protocol for 3D OOD
Here, we examine how external factors and hyperparameter choices influence OOD detection performance because these elements directly affect both the quality and reliability of the detection metrics. For instance, the characteristics of the dataset, such as point cloud density, class diversity, and the presence of noise, can alter how many true and false positives are identified. Similarly, hyperparameters like the distance threshold for matching predictions, the detection score cutoff, and the criteria used for sorting predictions determine which detections are considered valid. Even small adjustments in these parameters can lead to significant shifts in performance metrics, affecting the reproducibility and fairness of the evaluation. Therefore, we examine and compare these choices.
@@ IV-A Metrics and Matching
  Compute Euclidean distance between the M predictions (objects whose base detector prediction score exceeds $\delta_{\text{thresh}}$) and the NN GTs. Sort by order of confidence the distances, the boxes, and ”unknown” scores $p_{\text{OOD}}$;
@@
        if dist(pred, closest unmatched GT)<< $d_{\text{thresh}}$ then